OpenStudy (anonymous):

find the derivative of y=sin^-1(7x)

OpenStudy (anonymous):

\[\frac{ 7 }{ 1+x^2 }\]

OpenStudy (anonymous):

\[\frac{ d }{ dx }\sin^{-1}x=\frac{ 1 }{ 1+x^2 }\]

OpenStudy (anonymous):

and d/dx of 7x=7

OpenStudy (anonymous):

what is the derivative of arctan(x/3)

OpenStudy (anonymous):

do you use 1/ 1 + x^2

OpenStudy (anonymous):

post your new question in a new thread thanx

OpenStudy (anonymous):

sorry where do I do this

OpenStudy (anonymous):

close this question n post another 1